### v4

# FLOAT: Generative Motion Latent Flow Matching for Audio-driven Talking Portrait

Abstract
With the rapid advancement of diffusion-based generative models, portrait image animation has achieved remarkable results. However, it still faces challenges in temporally consistent video generation and fast sampling due to its iterative sampling nature. This paper presents FLOAT, an audio-driven talking portrait video generation method based on flow matching generative model. Instead of a pixel-based latent space, we take advantage of a learned orthogonal motion latent space, enabling efficient generation and editing of temporally consistent motion. To achieve this, we introduce a transformer-based vector field predictor with an effective frame-wise conditioning mechanism. Additionally, our method supports speech-driven emotion enhancement, enabling a natural incorporation of expressive motions. Extensive experiments demonstrate that our method outperforms state-of-the-art audio-driven talking portrait methods in terms of visual quality, motion fidelity, and efficiency.
Figure 1: FLOAT can generate a talking portrait video from a single source image and audio where the talking motion is generated by the motion latent flow matching. It can enhance the emotion-related talking motion by leveraging speech-driven emotion labels, a natural way of emotion-aware motion control.
$\star$\$$\star$\$footnotetext: This work was done during South Korea Mandatory Military Service at DeepBrain AI Inc.
1 Introduction
Animating a single image using a driving audio (i.e., audio-driven talking portrait generation) has gained significant attention in recent years for its great potential in avatar creation, video conferencing, virtual avatar chat, and user-friendly customer service. It aims to synthesize natural talking motion from audio signals, including accurate lip synchronization, rhythmical head movements, and fine-grained facial expressions. However, generating such motion solely from audio is extremely challenging due to its one-to-many correlation between audio and motion. In the earlier stage of this field, many works [58, 34, 23, 54, 9, 98] focus on generating accurate lip movements by relying on learned audio-lip alignment losses [10, 52].
To comprehensively extend the range of motion, some works [96, 52, 74] incorporate probabilistic generative models, such as VAE [35] and normalizing flow [60], turning the motion generation into probabilistic sampling. However, these models still lack expressiveness in generated motion due to the limited capacity of these generative models.
Recent talking portrait generation methods [76, 86, 8, 89, 31, 43, 51, 80, 25, 70], powered by diffusion-based generative models [27, 68], successfully mitigate this expressiveness issue. EMO [76] introduces a promising approach to this field [86, 80, 8, 89, 31] by employing a strong pre-trained image diffusion model (i.e., StableDiffusion [61]) and lifting it into video generation [29]. However, it still faces challenges in generating temporally coherent videos and achieving sampling efficiency, requiring tens of minutes for a few seconds of video. Moreover, they heavily rely on auxiliary facial prior, such as bounding boxes [76, 89], 2D landmarks and skeletons [94, 8, 31], or 3D meshes [86], which significantly restricts the diversity and the fidelity of head movements due to their strong spatial bias.
In this paper, we present FLOAT, an audio-driven talking portrait video generation model based on flow matching generative model in a motion latent space. Flow matching [42, 44] has emerged as a promising alternative to diffusion models due to its fast and high-quality sampling. By modeling talking motion within a learned motion latent space [85], we can more efficiently sample temporally consistent motion latents. This is achieved by a simple yet effective transformer-based [79] vector field predictor, inspired by DiT [55]. Since our motion latent space has orthogonal structure, our method can manipulate head motion of the generated video using its basis. Furthermore, our method supports natural emotion-aware motion enhancement driven by speech. Our contributions are summarized as follows:
  • We present, FLOAT, flow matching based audio-driven talking portrait generation model using a learned orthogonal motion latent space, enabling to generate talking portrait videos with reduced sampling steps.
  • We introduce a simple yet effective transformer-based flow vector field predictor for temporally consistent motion latent sampling, which also enables the speech-driven emotional controls.
  • Extensive experiments demonstrate that FLOAT achieves state-of-the-art performance compared to both diffusion- and non-diffusion-based methods.
2 Related Works
2.1 Diffusion Models and Flow Matching
Diffusion Models   Diffusion models or score-based generative models [61, 27, 67, 53, 68, 14] are generative models that gradually diffuse input signals into Gaussian noise and learn the denoising reverse process for the generative modeling. They have shown remarkable results in various generation tasks, such as unconditional image and video generation [55, 4, 18], text-to-image generation [61, 59, 62], text-to-video generation [4, 24], conditional image generation [94, 29], and 3D human generation [75, 71, 37].
Accelerating Diffusion Models   While diffusion models demonstrate superior performance, their iterative sampling nature still bottlenecks the efficient generation compared to VAEs [35], normalizing flow [60], and GANs [22]. To overcome this limitation, several works have been developed to boost the sampling speed of the diffusion models. StableDiffusion (SD) [61] partially mitigates this problem by moving the diffusion process from the pixel space to the spatial latent space, establishing itself as a pivotal framework among diffusion models. Another line of research has developed the sampling solvers [47, 48] based on ordinary differential equations (ODEs). Meanwhile, model distillation [26] has been introduced to transfer the knowledge of the learned diffusion models into a student model, enabling one (or a few) steps of generation [69, 49, 45, 32, 41]. However, these approaches involve substantial effort to create a well-trained diffusion model and suffer from training instability.
Flow Matching   Flow matching [42, 44] stands out as an alternative to diffusion models for its high sampling speed and competitive sample quality compared to diffusion models [42, 11, 39, 20, 57]. It belongs to the family of flow-based generative models, which estimates a transformation (referred to as a flow) between a prior distribution (e.g., Gaussian) and a target distribution. Unlike the normalizing flow [60, 15] that directly estimates the noise-to-data transformation under specific architectural constraints (e.g., affine coupling), flow matching regresses the time-dependent vector field that generates this flow by solving its corresponding ODEs [7] with flexible architectures. One specific design of flow matching is an optimal transport (OT) based one, which transforms the data distribution along the straight path with constant velocity [42].
Our audio-driven talking portrait method employs flow matching to generate the natural talking motions. Thanks to the architectural flexibility of flow matching, we use transformer-encoder architecture [79] to estimate the generating vector field, allowing us to take the video temporal consistency into account.
2.2 Audio-driven Portrait Animation
Audio-driven portrait animation is the task of generating a realistic talking portrait video using a single portrait image and driving audio [82, 100, 99, 96, 52]. Since audio-to-motion relation is basically a one-to-many problem, several works utilize additional facial prior for driving conditions, e.g., 2D facial landmarks [100, 80, 8, 31, 25, 86], 3D prior [50, 91, 96, 9, 51], or emotional labels [90, 30, 73]. In earlier stages, most works [58, 34, 23, 9] focused on generating accurate lip motion from audio by utilizing the lip-sync discriminator [10]. These approaches have advanced to generating audio-related head poses in a probabilistic way. For example, StyleTalker [52] uses normalizing flow [60, 15] to generate the head motion from audio, while SadTalker [96] uses audio-conditional variational inference [35] to learn the 3DMM coefficients [2], bridging the intermediate representations of a pre-trained portrait animator [83].
Meanwhile, several works [81, 30, 73, 87] focus on an emotion-aware talking portrait generation. In particular, EAMM [30] considers an emotion as the complementary displacement of facial motion, and learns these displacement from an emotion label extracted from the image.
Recent audio-driven talking portrait methods powered by diffusion models show remarkable results [51, 76, 89, 90, 31, 8, 80, 86, 43]. Specifically, EMO [76] and subsequent extensions [89, 8, 80, 86] utilize the pre-trained SD [61] as their backbone to leverage generative prior trained on the large-scale image datasets. They introduce additional modules, e.g., ReferenceNet [29] and Temporal Transformer [24], to preserve input identity and enhance the video temporal consistency, respectively. However, these modules introduces additional computational cost, requiring several minutes for a few seconds of video, and still suffer from video-level artifacts, such as noisy frames, and flickering.
VASA-1 [90] addresses the sampling time issue by sampling motion latents [16], producing lifelike talking portraits. Our method takes advantage of this approach. However, unlike [90], our motion latent space has a strong linear orthogonal structure represented by a computable basis, enabling to manipulate the generated motion at the test-time without external driving signals. Based on this orthogonality, we employ OT-based flow matching for motion latent sampling along a straight line with reduced sampling steps.
3 Preliminaries: (Conditional) Flow Matching
Figure 2: Overview of FLOAT. We encode the source image $S\in\mathbb{R}^{3\times H\times W}$ into the latent with the explicit identity-motion decomposition $w_{s}=w_{s\to r}+w_{r\to s}\in\mathbb{R}^{d}$. Given audio segments $a^{-L^{\prime}:L}\in\mathbb{R}^{(L^{\prime}+L)\times d_{a}}$ of the length $L^{\prime}+L$ and the reference motion $w_{r\to s}$in $\mathbb{R}^{d}$, and the speech-driven emotion label $w_{e}$in $\mathbb{R}^{7}$, a flow matching transformer estimates the generating vector field $v_{t}(\varphi_{t}(x_{0}),\mathbf{c}_{t};\theta)\in\mathbb{R}^{L\times d}$ from noisy motion latents, which is used to solve corresponding ODE and generates the motion latents $w_{r\to\hat{D}^{1:L}}$. Finally, the sequence of latents $w_{S\to\hat{D}^{1:L}}:=(w_{S\to r}+w_{r\to\hat{D}^{l}})_{l=1}^{L}$ are decoded into the video $\hat{D}^{1:L}\in\mathbb{R}^{L\times 3\times H\times W}$.
Let $x$in $\mathbb{R}^{d}$ be a data, $t$in [0,1][0,1][ 0 , 1 ] be the time, and $q$ be a unknown target distribution. We can define a flow as a time-dependent transformation $\varphi_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that transforms a tractable prior distribution $p_{0}$ to the distribution $p_{1}\approx q$. This flow $\varphi_{t}$ further introduces a probability flow path $p_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}_{>0}$ and a generating vector field $v_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ where $p_{t}$ is defined by the push-forwarding
$p_{t}(x)=p_{0}(\varphi_{t}^{-1}(x))\det\left|\frac{\partial\varphi_{t}^{-1}(x)}{\partial x}\right|,$ (1)
and $v_{t}$ generates $\varphi_{t}$ by means of an ordinary differential equation (ODE) [7]:
$\frac{d}{dt}\varphi_{t}(x)=v_{t}(\varphi_{t}(x))\quad\text{and}\quad\varphi_{0}(x)=x.$ (2)
Flow matching [42] aims to estimate the target generating vector field $u_{t}$ with a neural network parameterized by $\theta$:
$\mathcal{L}_{\text{FM}}(\theta):=\|v_{t}(x;\theta)-u_{t}(x)\|_{2}^{2},$ (3)
where $t$sim $\mathcal{U}[0,1]$ and $x$sim $p_{t}(x)$. However, the target generating vector field $u_{t}$ and the sample distribution $p_{t}$ are intractable. To address this issue, [42] proposes a method for constructing a “conditional" probability path $p_{t}(\cdot|x_{1})$ as well as target “conditional" vector field $u_{t}(\cdot|x_{1})$ using a sample $x_{1}$sim $q$ as a condition. And they prove that the following objective
$\mathcal{L}_{\text{CFM}}(\theta):=\|v_{t}(x;\theta)-u_{t}(x|x_{1})\|_{2}^{2},$ (4)
where $t$sim $\mathcal{U}[0,1]$ and $x$sim $p_{t}(x|x_{1})$, is equivalent to (3) with respect to the gradient $\nabla_{\theta}$.
One natural way of constructing $u_{t}(\cdot|x_{1})$ is a “straight line" that connects $x_{0}$sim $p_{0}$ and $x_{1}$sim $q$, drawing an optimal transport (OT) path with constant velocity [42]. Specifically, a linear time interpolation between $x_{0}$ and $x_{1}$ gives us the flow $x_{t}=\varphi_{t}(x)=(1-t)x_{0}+tx_{1}$, the conditional probability path $p_{t}(x|x_{1})$ defined via the affine transformation $p_{t}(x|x_{1})=\mathcal{N}(x|tx_{1},(1-t)^{2}I)$, and the target generating vector field $u_{t}(x|x_{1})=x_{1}-x_{0}$. This specific choice turns the objective (4) into
$\mathcal{L}_{\text{OT}}(\theta):=\|v_{t}((1-t)x_{0}+tx_{1};\theta)-(x_{1}-x_{0})\|_{2}^{2},$ (5)
where $t$sim $\mathcal{U}[0,1]$, $x_{0}$sim $p_{0}$, and $x_{1}$sim $q$, all of which are tractable.
Classifier-free Vector Field   [11] formulates a classifier-free vector field (CFV) technique for flow matching, which enables class-conditional sampling more controllable manner without any extra classifier trained on noisy trajectory. Formally, CFV compute the modified vector field $\tilde{v}_{t}$ by
$\tilde{v}_{t}(x_{t},c;\theta)\approx\gamma v_{t}(x_{t},c;\theta)+(1-\gamma)v_{t}(x_{t},c=\emptyset;\theta),$ (6)
where $\gamma$ denotes the guidance scale. $v_{t}(x_{t},c=\emptyset;\theta)$ is the predicted vector field without a driving condition $c$. For more details, please refer to [42, 11].
4 Method: Flow Matching for Audio-driven Talking Portrait
We provide an overview of FLOAT in Fig.˜2. Given source image $S$in $\mathbb{R}^{3\times H\times W}$, and a driving audio signal $a^{1:L}$in $\mathbb{R}^{L\times d_{a}}$ of length $L$, our method generates a video
$\hat{D}^{1:L}=(\hat{D}^{l})_{l=1}^{L}\in\mathbb{R}^{L\times 3\times H\times W}$ (7)
of $L$ frames, featuring audio-synchronized talking head motions, including both verbal and non-verbal motions. Our method consists of two phases. First, we pre-train a motion auto-encoder, which provides us with the expressive and smooth motion latent space for the talking portraits (Sec.˜4.1). Next, we employ OT-based flow matching [42] to generate a sequence of motion latents with a transformer-based vector field predictor using the driving audio, which is decoded to the talking portrait videos (Sec.˜4.2). We also incorporate speech-driven emotions as the driving conditions, achieving automatic emotion-aware talking portrait generation without any extra user input for emotion.
4.1 Motion Latent Auto-encoder
Recent talking portrait methods utilize the VAE of StableDiffusion (SD) [61] due to its rich semantic pixel-based latent space. However, they often struggle to generate temporally consistent frames when lifted to video generating tasks [76, 89, 8, 29, 101]. Thus, our first goal for realistic talking portrait is to obtain good motion latent space, capturing both global (e.g., head motion) and fine-grained local (e.g., facial expressions, mouth and pupil movement) dynamics.
Instead of VAE of SD, we pre-train a motion latent auto-encoder that encodes images into motion latents. This is done by training an auto-encoder to reconstruct a driving image from a source image sampled from the same video clip, enforcing the encoder to implicitly capture both temporally adjacent and distant motions. Inspired by LIA [85], we introduce a learned orthonormal basis that decomposes the motion along distinct orthogonal directions. Specifically, our motion auto-encoder encodes the source $S$ into the latent $w_{S}\in\mathbb{R}^{d}$ with following explicit decomposition:
$w_{S}:=w_{S\to r}+w_{r\to S},$ (8)
where $w_{S\to r}$in $\mathbb{R}^{d}$ is the identity latent and
$w_{r\to S}=\sum_{m=1}^{M}\lambda_{m}(S)\cdot\mathbf{v}_{m}\in\mathbb{R}^{d}$ (9)
is the motion latent with $\lambda(S)$ :=:=:= $\left(\lambda_{m}(S)\right)_{m=1}^{M}$in $\mathbb{R}^{M}$ being the source-dependent motion coefficients that span the learned source-agnostic motion basis $V:=\{\mathbf{v}_{m}\}_{m=1}^{M}\subseteq\mathbb{R}^{d}$. In this space, $\lambda_{m}(S)$ is the intensity of the motion direction $\mathbf{v}_{m}$.
As shown in Fig.˜6, our method enables motion editing of the sampled (generated) motion using only the basis $V$ and its orthogonality, as stated in Eq.˜15.
Improving Fidelity of Facial Components: $\mathcal{L}_{\text{comp-lp}}$     The expressiveness of generated motions and the image fidelity are determined by the motion space and the motion auto-encoder. However, as resolution increases, fine details in small facial regions (e.g., teeth, eyeballs) often get buried in large-scale dynamics. To address this issue, we propose a facial component perceptual loss $\mathcal{L}_{\text{comp-lp}}$ using [95, 66] that significantly improves the image fidelity (e.g., teeth and eyes) as well as fine-grained motions (e.g., eyeball and eyebrows movements). As shown in Fig.˜3, $\mathcal{L}_{\text{comp-lp}}$ allows us to generate high-fidelity facial components and their fine-grained motions without relying on pre-trained foundation models, such as StableDiffusion [61].
Figure 3: Efficacy of $\mathcal{L}_{\text{comp-lp}}$ for fine-grained motion and fidelity.
4.2 Flow Matching in Motion Latent Space
Armed with this linear orthogonal space, we employ OT-based flow matching [42, 44] for the motion sampling. Specifically, we predict a vector field $v_{t}(x_{t},\mathbf{c}_{t};\theta)$in $\mathbb{R}^{L\times d}$ where $x_{t}$ is the sample at flow time $t\in[0,1]$, and $\mathbf{c}_{t}$in $\mathbb{R}^{L\times h}$ represents the driving conditions for $L$ consequent frames. This vector field generates the flow $\varphi_{t}:[0,1]\times\mathbb{R}^{L\times d}\to\mathbb{R}^{L\times d}$ of $L$ frames by solving ODE (Eq.˜2). As illustrated in Fig.˜4, we build our vector field predictor upon the transformer encoder [79] architecture. Specifically, we adopt DiT [55] architecture, but decouple frame-wise conditioning from time-axis attention mechanism, which enables us to model temporally consistent motion latents.
In DiT [55], distinct semantic tokens are modulated by a single diffusion time step embedding and class embedding through adaptive layer normalization (AdaLN). In contrast, our vector field predictor modulates each $l$-th input latent with its corresponding $l$-th condition and then combines their temporal relations through a masked self-attention layer that attends to $2\cdot T$ neighboring frames. Formally, for each $l$-th frame, frame-wise AdaLN and frame-wise gating are computed by
$\gamma_{i}^{l}\times\text{LN}(X_{t}^{l})+\beta_{i}^{l}\in\mathbb{R}^{h}\quad\text{and}\quad\alpha_{i}^{l}\times X_{t}^{l}\in\mathbb{R}^{h},$ (10)
[…]
Speech-driven Emotion Enhancement   How can we make talking motions more expressive and natural? During talking, humans naturally reflect their emotions through their voices, and these emotions influence talking motions. For instance, a person who speaks sadly may be more likely to shake the head and avoid eye contact. This non-verbal motion derived from emotions crucially impacts the naturalness of a talking portrait.
Existing works [81, 30, 90] use image-emotion paired data or image-driven emotion predictor [63] to generate the emotion-aware motion. In contrast, we incorporate speech-driven emotions, a more intuitive way of controlling emotion for audio-driven talking portrait. Specifically, we utilize a pre-trained speech emotion predictor [56] that produces softmax probabilities of seven distinct emotions: angry, disgust, fear, happy, neutral, sad, and surprise, which we then input into the vector field predictor.
However, as people do not always speak with a single, clear emotion, determining emotions solely from audio is often ambiguous [30]. Naive introduction of speech-driven emotion can make emotion-aware motion generation more challenging. To address this issue, we inject the emotions together with other driving conditions at training phase and modify them at inference phase.
Figure 4: Frame-wise vector field predictor block at inference.
Driving Conditions   We concatenate the audio representation $a^{1:L}$in $\mathbb{R}^{L\times d_{a}}$ of a pre-trained Wav2Vec2.0 [1], the speech emotion label $w_{e}$in $\mathbb{R}^{7}$, and the source motion latent $w_{r\to S}$in $\mathbb{R}^{d}$. Next, we add the flow time step embedding $\text{Emb}(t)$in $\mathbb{R}^{h}$ to these conditions, producing $\mathbf{c}_{t}$in $\mathbb{R}^{L\times h}$ via a linear layer, ToCondition, as depicted in Fig.˜2, where $\text{Emb}(t)$ is computed using the sinusoidal position embedding [79].
Training   We train FLOAT by reconstructing a target vector field computed from driving frames using the corresponding audio segments and a source motion latent. We choose a pair of driving motions and corresponding audio ($w_{r\to D^{1:L}}$, $a^{1:L}$), and construct the target vector field $u_{t}(x|w_{r\to D^{1:L}})=w_{r\to D^{1:L}}-x_{0}$in $\mathbb{R}^{L\times d}$ with noisy input $\varphi_{t}(x_{0})=(1-t)x_{0}+tw_{r\to D^{1:L}}$ ($t$sim $\mathcal{U}[0,1]$ and $x_{0}$sim $\mathcal{N}(0^{1:L},I)$).
For smooth transitions of sequences longer than the window length $L$, we incorporate last $L^{\prime}$ audio features and motion latents $w_{r\to D^{-L^{\prime}:0}}$ from the preceding window as additional input.
The flow matching objective $\mathcal{L}_{\text{OT}}(\theta)$ is defined by
$\mathcal{L}_{\text{OT}}(\theta)$ $=\|v_{t}^{1:L}(x_{t},\mathbf{c}_{t};\theta)-u_{t}(x|w_{r\to D^{1:L}})\|,$ (11)
$+\|v_{t}^{-L^{\prime}:0}(x_{t},\mathbf{c}_{t};\theta)-w_{r\to D^{-L^{\prime}:0}}\|,$
where $x_{t}$ :=:=:= $[w_{r\to D^{-L^{\prime}:0}}|~\varphi_{t}(x_{0})]$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$ is the concatenated input, $\mathbf{c}_{t}$in $\mathbb{R}^{(-L^{\prime}+L)\times h}$ is the driving condition consisting of $[t,w_{r\to S},w_{e},a^{1:L},a^{-L^{\prime}:0}]$. Note that $w_{e}$ and $w_{r\to S}$ are shared across the $L^{\prime}+L$ frames. We incorporate a velocity loss [75] to supervise temporal consistency:
$\mathcal{L}_{\text{vel}}(\theta)=\|\Delta v_{t}-\Delta u_{t}\|,$ (12)
where $\Delta v_{t}$ and $\Delta u_{t}$ are the one-frame difference along the time-axis for the prediction $v_{t}$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$ and the target $[w_{r\to D^{-L^{\prime}:0}}|~u_{t}]$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$, respectively.
The total objective $\mathcal{L}_{\text{total}}(\theta)$ is
$\mathcal{L}_{\text{total}}(\theta)=\lambda_{\text{OT}}\mathcal{L}_{\text{OT}}(\theta)+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}(\theta),$ (13)
where $\lambda_{\text{OT}}$ and $\lambda_{\text{vel}}$ are the balancing coefficients. During training, we apply dropout to $w_{r}$, $w_{e}$, and $a^{1:L}$ with a probability of 0.10.10.1 for CFV. Additionally, we apply dropout to the preceding audio and motion latents with a probability 0.50.50.5 for smooth transition in the initial window.
Table 1: Quantitative comparison results with state-of-the-art methods on HDTF [97] / RAVDESS [46]. The best result for each metric is in bold, and the second-best result is underlined. : evaluated with raw $\times 256$ 256 resolution outputs.
Method Image & Video Generation Lip Synchronization
FID $\downarrow$ FVD $\downarrow$ CSIM $\uparrow$ E-FID $\downarrow$ P-FID $\downarrow$ LSE-D $\downarrow$ LSE-C $\uparrow$
SadTalker [96] 71.952 / 119.430 339.058 / 376.294 0.644 / 0.644 1.914 / 3.500 1.456 / 2.045 7.947 / 7.273 7.305 / 4.748
EDTalk [74] 50.078 / 75.020 211.284 / 304.933 0.626 / 0.676 1.579 / 3.468 0.054 / 0.090 8.123 / 7.682 7.623 / 5.318
AniTalker [43] 39.512 / 70.430 184.454 / 265.341 0.643 / 0.725 1.830 / 2.330 0.092 / 0.126 7.907 / 8.176 7.288 / 4.555
Hallo [89] 25.363 / 57.648 197.196 / 375.557 0.869 / 0.860 1.039 / 2.492 0.037 / 0.050 7.792 / 7.613 7.582 / 4.795
EchoMimic [8] 33.552 / 81.839 296.757 / 320.220 0.823 / 0.805 1.234 / 3.201 0.023 / 0.047 8.903 / 8.161 6.242 / 4.144
FLOAT (Ours) 21.100 / 31.681 162.052 / 166.359 0.843 / 0.810 1.229 / 1.367 0.032 / 0.031 7.290 / 6.994 8.222 / 5.730
Figure 5: Qualitative comparison results with state-of-the-art methods on HDTF [97] / RAVDESS [46]. Please refer to supplementary videos. Note that we additionally provide a video comparison with EMO [76] and VASA-1 [90] using their video demonstration.
Inference   During inference, we sample the generating vector field from noise $x_{0}$, using the driving conditions $w_{r\to S}$, $w_{e}$, and $a^{1:L}$, as well as the $L^{\prime}$ frames of preceding audio and generated motion latents.
We extend the CFV [11] to an incremental CFV to separately adjust the audio and emotion, inspired by [3]:
$\tilde{v}_{t}$ $\approx v_{t}(x_{0},\mathbf{c}_{t}|_{\{a^{1:L},w_{e}\}})$
$+\gamma_{a}\left[v_{t}(x_{0},\mathbf{c}_{t}|_{w_{e}})-v_{t}(x_{0},\mathbf{c}_{t}|_{\{a^{1:L},w_{e}\}}\right]$
$+\gamma_{e}\left[v_{t}(x_{0},\mathbf{c}_{t})-v_{t}(x_{0},\mathbf{c}_{t}|_{w_{e}})\right],$ (14)
where $\gamma_{a}$ and $\gamma_{e}$ are the guidance scales for audio and emotion, respectively. $\mathbf{c}_{t}|_{\{x,y\}}$ denotes the driving condition without the condition $x$ and $y$. We set $\gamma_{a}=2$ and $\gamma_{e}=1$ based on the ablation studies on $\gamma_{a}$ and $\gamma_{e}$ provided in supplementary materials.
After sampling, ODE solver receives the estimated vector field to compute the motion latents through numerical integration. We empirically find that FLOAT can generate reasonable motion with around 10 number of function evaluations (NFE). Please refer to supplementary videos.
Lastly, we add the source identity latent to the generated motion latents and decode them into video frames using the motion latent decoder.
5 Experiments
5.1 Dataset and Pre-processing
For training the motion latent auto-encoder, we use three open-source datasets: HDTF [97], RAVDESS [46], and VFHQ [88]. When training FLOAT, we exclude VFHQ because it does not support the synchronized audio. HDTF [97] is for high-definition talking face generation, containing videos of over 300 unique identities. RAVDESS [46] includes more than 2,400 emotion-intensive videos of 24 different identities. VFHQ [88] is designed for high-resolution video super-resolution and includes a large number of unique identities, which compensates the limited number of identities of the preceding datasets. Following the strategy of [65], we first convert each video to 25 FPS and resample the audio into 16 kHz. Then, we crop and resize the facial region to $512^{2}$ resolution [5]. After the pre-processing, for HDTF, we use a total of 11.3 hours of 240 videos featuring 230 different identities for training, and videos of 78 different identities, each 15 seconds long, for test. For RAVDESS, we use videos of 22 identities for training, and videos of the remaining 2 identities for test, with each 3-4 seconds long and representing 14 emotional intensities. Note that the identities in the training and test are disjoint in both datasets.
5.2 Implementation Details
[…]
5.3 Evaluation
Metrics and Baselines   For evaluating the image and video generation quality, we measure Fréchet Inecption Distance (FID) [64] and 16 frames Fréchet Video Distance (FVD) [78]. For facial identity, expression and head motion, we measure Cosine Similarity of identity embedding (CSIM) [12], Expression FID (E-FID) [76] and Pose FID (P-FID), respectively. Lastly, we measure Lip-Sync Error Distance and Confidence (LSE-D and LSE-C [58]) for audio-visual alignment.
We compare our method with state-of-the-art audio-driven talking portrait methods whose official implementations are publicly available. For non-diffusion methods, we compare with SadTalker [96] and EDTalk [74]. For diffusion methods, we compare with AniTalker [43], Hallo [89], and EchoMimic [8].
Comparison Results   In Tab.˜1 and Fig.˜5, we show the quantitative and qualitative comparison results, respectively. FLOAT outperforms other methods on most of the metrics and visual quality in both datasets.
Additionally, we provide video comparison results with EMO [76] and VASA-1 [90] in the supplementary materials, using their demonstration videos due to the infeasibility of direct implementation.
Figure 6: Test-time pose editing using $\lambda$-control ($\lambda_{15}(\hat{D})\pm 10$).
5.4 Applications
Test-time Pose Editing via Orthonormal Basis $V$   Since FLOAT learns the underlying motion latent structure, it is natural to assume that for any sampled motion latent $w_{r\to\hat{D}}$, there exist motion coefficients $\{\lambda_{m}(\hat{D})\}_{m=1}^{M}$ satisfying the representation in Eq.˜9: $w_{r\to\hat{D}}$ $=$ $\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m}$.
We can always compute these coefficients in closed form by taking inner products between the sampled motion $w_{r\to\hat{D}}$ and the learned orthonormal basis $V$:
$\langle w_{r\to\hat{D}},~\mathbf{v}_{k}\rangle=\langle\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m},~\mathbf{v}_{k}\rangle=\lambda_{k}(\hat{D}),$ (15)
where $\langle\mathbf{v}_{m},\mathbf{v}_{k}\rangle=\delta_{m,k}$ and $\delta$ is Kronecker delta. At this point, we can edit the sampled motions by editing the corresponding coefficients (e.g., via linear operation) and combining them back into the motion latent. As shown in Fig.˜6, it allows us to control head direction without interfering with other motions due to the orthogonality of the basis. We refer to this test-time editing technique as $\lambda$-control.
Additional Driving Signals   In Fig.˜7 and Tab.˜2, we experiment with additional driving conditions, head poses and image-driven emotion labels, to explore additional controllability in our method. We employ 3DMM head pose parameters $p$in $\mathbb{R}^{6}$ [2] extracted by [13]. We concatenate a sequence of pose parameters $p^{1:L}$in $\mathbb{R}^{L\times 6}$ with the other driving conditions, and then map them to $c_{t}^{1:L}$in $\mathbb{R}^{L\times h}$. We also experiment on image-driven emotion [63] for frame-wise emotion control rather than the long-term emotion enhancement. FLOAT can effectively accommodate these additional conditions, highlighting its flexibility across diverse control signals.
Redirecting Speech-driven Emotion  Since FLOAT learns diverse emotions in the emotion-intensive data distribution [46], the generated emotion-aware motion can be modified by redirecting the speech-driven emotion label toward a different emotion at inference time. As illustrated in Fig.˜8, this technique is particularly beneficial for manual redirection when the emotion predicted from speech is complex or ambiguous.
Figure 7: Additional conditioning results of FLOAT. 3DPose, S2E, and I2E denote 3D head pose parameters [13], speech-to-emotion [56], and image-to-emotion [63], respectively.
Table 2: Quantitative results of FLOAT with additional conditions (HDTF [97] / RAVDESS [46]). S2E, I2E, and 3DPose denote speech-to-emotion [56], image-to-emotion [63], and 3DMM pose parameters [13], respectively.
Configurations FID $\downarrow$ FVD $\downarrow$ E-FID $\downarrow$ P-FID $\downarrow$ LSE-D $\downarrow$
A FLOAT (Ours) 21.100 / 31.681 162.052 / 166.359 1.229 / 1.367 0.032 / 0.031 7.290 / 6.994
B A + 3DPose 19.721 / 29.721 126.663 / 112.894 0.926 / 1.152 0.012 / 0.016 7.516 / 7.047
C A - S2E 21.235 / 32.035 155.032 / 166.866 1.254 / 1.502 0.031 / 0.025 7.264 / 7.222
D A - S2E + I2E 21/528 / 31.609 158.577 / 162.369 1.158 / 1.305 0.034 / 0.022 7.183 / 7.150
5.5 Ablation Studies
Ablation on Frame-wise AdaLN   We compare frame-wise AdaLN (and gating) followed by masked self-attention to separate conditioning from attending, with a cross-attention that performs conditioning and attending simultaneously. As shown in Tab.˜3, both approaches achieve competitive image and video quality, while frame-wise AdaLN provides better expression generation and lip synchronization. We observe that frame-wise AdaLN can achieve more diverse head motions than the cross-attention. Please refer to supplementary videos.
Figure 8: Redirecting the unclear emotion prediction to a desirable one-hot encoding, which can be further intensified by the CFV.
Ablation on Flow Matching   We compare flow matching with two types of diffusion models: $\epsilon$-prediction (noise) and $x_{0}$-prediction (signal) [59, 75]. In both cases, we adopt our vector predictor architecture as denoising networks. We adopt diffusion training settings of VASA-1 [90] (500 diffusion steps with a cosine noise scheduler [53] and 50 DDIM denoising steps) for the indirect comparison with [90]. Notably, diffusion and flow matching achieve competitive results on image quality while the latter achieves the better lip synchronization. In Fig.˜9, we compare the forward pass efficiency by measuring frames per second (FPS) of each model. Thanks to the compact motion latent representation and OT-based flow matching, FLOAT achieves the highest FPS, superior lip-sync performance, dynamic head motion, and the lowest NFEs.
Table 3: Ablation studies of FLOAT on HDTF [97]. The best result for each metric is in bold, and the second-best result is underlined.
Method FID $\downarrow$ FVD $\downarrow$ E-FID $\downarrow$ LSE-D $\downarrow$ # NFEs $\downarrow$
Ours (w. Cross-Attn.) 21.873 162.702 1.452 7.757 10
Ours (w. Diff., $\epsilon$-pred.) 21.190 161.666 1.213 9.922 50
Ours (w. Diff., $x_{0}$-pred.) 21.697 162.847 1.278 9.048 50
FLOAT (Ours) 21.100 162.052 1.229 7.290 10
Figure 9: Comparison of the forward pass efficiency. We compute FPS on a single NVIDIA V100 GPU.
6 Conclusion
We proposed FLOAT, a flow matching based audio-driven talking portrait generation model leveraging a learned motion latent space. We introduced a transformer-based vector field predictor, enabling temporally consistent motion generation. Additionally, we incorporated speech-driven emotion labels into the motion sampling process to improve the naturalness of the audio-driven talking motions. FLOAT addresses current core limitations of diffusion-based talking portrait video generation methods by reducing the sampling time through flow matching while achieving the remarkable sample quality. Extensive experiments verified that FLOAT achieves state-of-the-art performance in terms of visual quality, motion fidelity, and efficiency.
Discussion   We leave further discussion considering limitations, future work, and ethical considerations in the supplementary materials.
Appendix A More on Motion Latent Auto-encoder
In this section, we provide more details on our motion latent auto-encoder, including its model architecture, dataset, and training strategy.
A.1 Model
We provide a detailed model architecture of our motion latent auto-encoder in Fig.˜17.
In Fig.˜11(a), Fig.˜11(b), Fig.˜11(c), and Fig.˜11(d), we present visualization results of the latent decomposition
$w_{S}=w_{S\to r}+w_{r\to S}\in\mathbb{R}^{d}$ (16)
of a source image $S$, following the approach of [85]. Notably, the identity latent $w_{r\to S}$ is decoded into image featuring the average head pose, expression, and field of view in pixel space.
Figure 10: Ablation study on Facial Component Loss $\mathcal{L}_{\text{comp-lp}}$. It significantly improves the image fidelity of facial component (e.g., teeth, highlighted in red box) and fined-grained motion (eyeball movement, highlighted in yellow box).
Table 4: Quantitative comparison result (Same-identity) of motion latent auto-encoders on HDTF [97] / RAVDESS [46] / VFHQ [88]. The best result for each metric is in bold. : Results generated by official implementation ($\times 256$ 256)
Method FID $\downarrow$ FVD $\downarrow$ LPIPS $\downarrow$ E-FID $\downarrow$ P-FID $\downarrow$
LIA [85] 47.481 / 67.541 / 89.209 172.195 / 130.836 / 342.964 0.184 / 0.122 / 0.245 1.279 / 1.153 / 1.106 0.120 / 0.005 / 0.013
Ours (w.o. $\mathcal{L}_{comp-lp}$) 21.061 / 28.866 / 46.950 150.340 / 103.145 / 299.757 0.110 / 0.072 / 0.165 1.369 / 1.157 / 0.872 0.011 / 0.010 / 0.014
Ours 19.803 / 23.350 / 43.992 147.089 / 100.345 / 291.560 0.108 / 0.062 / 0.161 1.334 / 1.053 / 1.006 0.010 / 0.008 / 0.012
(a) Source, $S$
(b) Driving, $D$
(c) Identity, $w_{S\to r}$
(d) Reconstruction, $\hat{D}$
(e) Component mask
(f) Component diff
Figure 11: Visualization results of the motion latent auto-encoder.
A.2 Improving Fidelity of Facial Components
Facial Components: Texture vs. Structure   As highlighted in face restoration work [84], facial components such as eyeballs and teeth play a important role in the perceptual quality of generated images. It treats the issue as a lack of texture (lying in high frequencies) and mitigate it by introducing facial component discriminators with the gram matrix statistics matching. This approach is appropriate in face restoration, where training objective is to reconstruct a clear image from a degraded one that maintains the same spatial structure, ensuring that the low-frequency structure preserved.
However, in the context of training a motion auto-encoder, spatial mismatches are inevitably involved. Therefore, naively applying such discriminators proves ineffective. Instead, achieving high-fidelity facial components in a motion auto-encoder is more closely related to structural problems (lying in low frequencies) than to texture issues as shown in Fig.˜11(f).
Facial Component Perceptual Loss $\mathcal{L}_{\text{comp-lp}}$   We introduce a simple yet effective facial component perceptual loss, which leverages the standard perceptual loss $\mathcal{L}_{\text{lp}}$ [95] known for its ability to capture structural features lying in low frequencies. Formally, the facial component perceptual loss is defined by
$\sum_{i=1}^{N}\frac{1}{|M_{i}|}\|M_{i}\otimes\phi_{i}(\hat{D})-M_{i}\otimes\phi_{i}(D)\|_{1},$ (17)
where $D$ is the driving, $\hat{D}$ is the generated image, $N$ is the number of feature pyramid scales, $\phi_{i}(X)$ is the $i$-th feature of the input image $X$ computed by VGG-19 [66, 95], $M_{i}$ is the binary mask of the facial components that has same size with $\phi_{i}(X)$, and $|M_{i}|$ is the sum of all values in the binary mask $M_{i}$. We adopt a single perceptual loss with $N=4$ scales of VGG-19 feature pyramids. It is worth noting that we mask all the multi-resolution features (not only the image).
To compute the facial component mask $M_{i}$, we utilize an off-the-shelf face segmentation model [92] for tight mouth regions and face landmark detector [5] for the bounding box regions of the eyes as illustrated in Fig.˜11(e).
In Tab.˜4, we conduct ablation studies on motion latent auto-encoders. Notably, $\mathcal{L}_{\text{comp-lp}}$ is consistently improves the image fidelity over three datasets. As illustrated in Fig.˜10, an additional advantage of $\mathcal{L}_{\text{comp-lp}}$ is its ability to directly supervise fine-grained motion (often neglected due to large head motion) such as eyeball movement without any external driving conditions such as eye-gazing direction [17].
A.3 Training Objective
We train our motion latent auto-encoder by reconstructing a driving image $D$ from a source image $S$, both sampled from the same video clip.
The total loss function $\mathcal{L}_{\text{total}}$ for the motion latent auto-encoder is defined as
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{L1}+\lambda_{\text{lp}}\mathcal{L}_{\text{lp}}+\lambda_{\text{comp-lp}}\mathcal{L}_{\text{comp-lp}}$
$+\lambda_{\text{full-adv}}\mathcal{L}_{\text{full-adv}}$
$+\lambda_{\text{eye-adv}}\mathcal{L}_{\text{eye-adv}}+\lambda_{\text{eye-FSM}}\mathcal{L}_{\text{eye-FSM}}$
$+\lambda_{\text{lip-adv}}\mathcal{L}_{\text{lip-adv}}+\lambda_{\text{lip-FSM}}\mathcal{L}_{\text{lip-FSM}},$ (18)
where $\lambda_{\text{lp}}$, $\lambda_{\text{comp-lp}}$, $\lambda_{\text{eye-adv}}$, $\lambda_{\text{eye-FSM}}$, $\lambda_{\text{lip-adv}}$, $\lambda_{\text{lip-FSM}}$, and $\lambda_{\text{full-adv}}$ are the balancing coefficients. Here, $\mathcal{L}_{L1}$ is the L1 loss, and $\mathcal{L}_{\text{lp}}$ is the VGG-19 [66] based multi-scale perceptual loss [95] similar to $\mathcal{L}_{\text{comp-lp}}$. We incorporate 2-scale discriminator $\mathcal{L}_{\text{full-adv}}$ with the non-saturating loss:
$\mathcal{L}_{\text{full-adv}}=-\log[\text{Disc}_{\text{full}}(\hat{D})],$ (19)
where Disc denotes a discriminator adopted from [33]. To improve the fidelity of the facial components, we also incorporate the facial component discriminators with the feature style matching (FSM) [84],
$\mathcal{L}_{x\text{-adv}}$ $=-\log[\text{Disc}_{x}(\hat{D}_{x})],$ (20)
$\mathcal{L}_{x\text{-FSM}}$ $=\|\text{Gram}(\psi(D_{x}))-\text{Gram}(\psi(\hat{D}_{x}))\|_{1},$ (21)
where $x\in\{\text{eye},\text{lip}\}$. $D_{x}$ and $\hat{D}_{x}$ represent the region of interest (RoI) for the component $x$ in the driving $D$ and reconstruction $\hat{D}$, respectively. Gram is a gram matrix calculation [21] and $\psi$ is the multi-resolution features extracted by the learned component discriminators.
Figure 12: Comparison results with EMO [76] and VASA-1 [90] based on their demonstration videos. Please note that their implementation are unavailable.
A.4 Implementation Details
We set the balancing coefficients $\lambda_{\text{lp}}=10$, $\lambda_{\text{comp-lp}}=100$, $\lambda_{\text{eye-adv}}=1$, $\lambda_{\text{eye-FSM}}=100$, $\lambda_{\text{lip-adv}}=1$, $\lambda_{\text{lip-FSM}}=100$, and $\lambda_{\text{full-adv}}=1$. We employ Adam optimizer [36] with a batch size of 8 and a learning rate of $2\cdot 10^{-4}$. Entire training takes about 9 days for 460k steps on a single NVIDIA A100 GPU.
For training our motion latent auto-encoder, we use VFHQ [88] to supplement the limited number of identities provided by HDTF [97] and RAVDESS [46]. After the same pre-processing, remaining 14,362 video clips are used for training, and 49 video clips are used for test, respectively.
Appendix B More on FLOAT
In this section, we provide more details on FLOAT, including model, experiments, and further results.
In Fig.˜18, we provide a detailed model architecture for the driving conditions $\mathbf{c}_{t}$.
B.1 Evaluation Metrics
We provide further details of following metrics.
  • LPIPS [95] is used to measure the perceptual similarity between reconstructed image and real image based on the pre-trained AlexNet features [38].
  • FID [64] aims to measure the distance between the feature distributions of real and generated datasets. It is computed as:
    $\|\mu_{r}-\mu_{g}\|_{2}^{2}+\text{Tr}(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{\frac{1}{2}}),$ (22)
    where $\mu_{r}$, $\Sigma_{r}$ and $\mu_{g}$, $\Sigma_{g}$ are the means and covariances of the pre-trained InceptionNet [72] features from the real and generated datasets, respectively.
  • FVD [78] is a variant of FID [64], which is used to measure the spatio-temporal consistency between the real and generated datasets by leveraging the features of pre-trained video model [6]. We compute this using 16 frames with a sliding window manner for each video.
  • CSIM [12] measures face similarity between the two face images by computing the cosine similarity between the pre-trained ArcFace features [12] of two images.
  • E-FID [76] aims to measure expression similarity by computing the FID score (Eq.˜22) of 3DMM expression parameters (64-dim) [13] of generated videos and real videos.
  • P-FID aims to measure the head pose similarity by computing the FID score (Eq.˜22) of 3DMM pose parameters (6-dim) [13] of generated videos and real videos.
  • LSE-D and LSE-C [58] measure lip synchronization using the pre-trained SynNet [10]. LSE-D computes the distance between the predicted audio embedding and the predicted video embedding, while LSE-C represents the confidence of synchronization.
B.2 Baselines
For non-diffusion-based methods, we compare with SadTalker [96] and EDTalk [74]. For diffusion-based methods, we compare with AniTalker [43], Hallo [89], and EchoMimic [8].
  • SadTalker [96] employs an audio-conditional variational auto-encoder (VAE) to synthesize the head motion and eye blink in a probabilistic way.
  • EDTalk [74] uses normalizing for audio-driven head motion generation and can separately control the lip and head motion.
  • AniTalker [43] introduces a diffusion model to the learned motion latent space (similar to FLOAT) along with a variance adapter to improve the motion diversity. We use HuBERT audio feature-based implementation [28] for improved lip synchronization and apply default guidance scales and denoising steps of the official implementation.
  • Hallo [89] uilizes the pre-trained StableDiffusion [61] as its image generator, incorporating a hierarchical audio attention module to separately control lip synchronization, expression, and head pose. We use default guidance scales and denoising steps provided in the official implementation.
  • EchoMimic [8] is also StableDiffusion-based method, which leverages facial skeleton as additional driving signals. We use the default guidance scales and denoising steps provided in the official implementation.
  • It is worth noting that we compare with two superior works EMO [76] and VASA-1 [90] based on their demonstration videos due to their unavailable implementation. We highly recommend referring to ‘01_\__EMO_\__VASA-1_\__Comparison///xxxx.mp4’.
B.3 More on Experiments
For evaluating our method, we use the first frame of each video clip as the source image. We use the first-order Euler method [42] as our ODE solver. We experimentally find that other ODE solvers, such as mid-point and Dopri5, do not lead to significant performance improvements.
Table 5: Ablation studies of the different NFE of ODE on HDTF [97]. FPS is computed on a single NVIDIA V100 GPU.
Ours-NFE FID $\downarrow$ FVD $\downarrow$ E-FID $\downarrow$ LSE-D $\downarrow$ FPS $\uparrow$
Ours-2 21.785 178.831 1.542 7.559 45.22
Ours-5 21.440 164.463 1.331 7.155 44.74
Ours-10 (default) 21.100 162.052 1.229 7.290 41.37
Ours-20 21.158 164.392 1.293 7.343 38.20
Ablation on NFE   In general, increasing the number of function evaluation (NFE) reduces the solution error of ODEs. As shown in Tab.˜5, even with small NFE =2=2= 2, FLOAT can achieve competitive image quality (FID) and lip synchronization (LSE-D). However, it struggles to capture consistent and expressive motions (FVD and E-FID), resulting in shaky head motion and a static expression. This is because FLOAT generates the motion in the latent space, while image fidelity is determined by the auto-encoder. We provide supplementary videos, illustrating the impact of different NFE (Number of Function Evaluations). Notably, with a small NFE of 2, the generated images exhibit good quality, but the head movements appear temporally unstable, and emotions may be exaggerated. Please refer to supplementary videos for temporal jitters of low NFE.
Table 6: Ablation studies of the audio guidance scale $\gamma_{a}$ and the emotion guidance scale $\gamma_{e}$ on RAVDESS [46].
Guidance scales FID $\downarrow$ FVD $\downarrow$ E-FID $\downarrow$ LSE-D $\downarrow$
$\gamma_{a}$=1,   $\gamma_{e}$=1 33.066 171.047 1.555 7.049
$\gamma_{a}$=1,   $\gamma_{e}$=2 31.844 166.041 1.334 7.212
$\gamma_{a}$=2,   $\gamma_{e}$=1  (default) 31.681 166.359 1.367 6.994
$\gamma_{a}$=2,   $\gamma_{e}$=2 32.253 162.658 1.351 6.994
Ablation on Guidance Scales   In Tab.˜6, we conduct ablation studies on guidance scales: $\gamma_{a}$ and $\gamma_{e}$, with the emotion intensive dataset RAVDESS [46]. Note that increasing $\gamma_{a}$ leads to better temporal consistency (FVD) and lip synchronization quality (LSE-D). Moreover, increasing $\gamma_{e}$ improves video consistency (FVD) and expressiveness (E-FID). This enables balanced control over emotional audio-driven talking portrait generation.
In Fig.˜20, we visualize the effect of different emotion guidance scale $\gamma_{e}$. For this experiments, the predicted speech-to-emotion label is disgust with $\%9$ probability. Notably, as increasing $\gamma_{e}$ from 0 to 2, we can observe that emotion-related expressions and motions are enhanced.
Figure 13: Ablation results on frame-wise AdaLN and flow matching. Please refer to supplementary video for notable differences.
[…]
To compare against flow matching, we implement two diffusion models with distinct parameterizations: $\epsilon$-prediction and $x_{0}$-prediction. For $\epsilon$-prediction, we directly predict Gaussian noise by the noise predictor $s(\cdot;\theta)$ parameterized by $\theta$ with the following simple loss:
$\mathcal{L}_{\text{simple, noise}}(\theta)=\|s(x_{t},\mathbf{c}_{t};\theta)-\epsilon\|_{2}^{2},$ (23)
where $t$sim $\mathcal{U}[0,1]$, $\epsilon$sim $\mathcal{N}(0^{-L^{\prime}:L},I)$, and the noise input $x_{t}$in $\mathbb{R}^{(L^{\prime}+L)\times d}$ is sampled from a forward diffusion process $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ [27]. In our case, $x_{t}$ is noisy motion latents at diffusion time step $t$, starting from $t=0$ with $x_{0}$ $=$ $w_{r\to D^{1:L}}$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$.
For $x_{0}$-prediction, we predict a clean sample $x_{0}$, instead of noise [59], by the predictor $s(\cdot;\theta)$ with the following simple loss:
$\mathcal{L}_{\text{simple},x_{0}}(\theta)=\|s(x_{t},\mathbf{c}_{t};\theta)-x_{0}\|_{2}^{2}.$ (24)
We also incorporate a velocity loss [75]:
$\mathcal{L}_{\text{vel},x_{0}}(\theta)=\|\Delta s-\Delta x_{0}\|_{2}^{2},$ (25)
where $\Delta s$ and $\Delta x_{0}$ are the one-frame difference along the time-axis for $s$ and $x_{0}$, respectively. The total loss $\mathcal{L}_{\text{total},x_{0}}(\theta)$ is
$\mathcal{L}_{\text{total},x_{0}}(\theta)=\mathcal{L}_{\text{simple},x_{0}}(\theta)+\mathcal{L}_{\text{vel},x_{0}}(\theta).$ (26)
For reverse process, we use the DDIM [67] sampler with 50 denoising steps.
In our implementation, both $\epsilon$-prediction and $x_{0}$-prediction achieve the best results with guidance scales $\gamma_{a}=\gamma_{e}=1$ (default). In Fig.˜13, Fig.˜21 and Fig.˜22, we provide qualitative comparisons between these approaches and FLOAT. Notably, the cross-attention exhibits less diverse head motions compared to FLOAT, while diffusion-based approaches struggle to generate temporally stable lip and head motion, often resulting in out-of-sync movements or motion artifacts.
Appendix C Additional Results
C.1 Additional Comparison Results
We provide additional comparison results with baselines in Fig.˜24, Fig.˜25, and Fig.˜26.
C.2 Out-of-distribution (OOD) Results
In Fig.˜19 and Fig.˜20, we present additional out-of-distribution results, including paintings, non-English speech, and singing.
C.3 User Study
[…]
Method
Lip Sync
Accuracy
Natural
Head Motion
Teeth
Clarity
Natural
Emotion
Overall
Visual Quality
SadTalker [96] 2.20 $\pm$ 0.35 2.03 $\pm$ 0.26 1.53 $\pm$ 0.19 1.80 $\pm$ 0.28 1.97 $\pm$ 0.23
EdTalk [74] 2.50 $\pm$ 0.34 2.60 $\pm$ 0.28 1.17 $\pm$ 0.17 2.07 $\pm$ 0.36 1.83 $\pm$ 0.27
AniTalker [43] 2.70 $\pm$ 0.31 3.00 $\pm$ 0.30 2.13 $\pm$ 0.27 3.17 $\pm$ 0.27 2.63 $\pm$ 0.26
Hallo [89] 3.30 $\pm$ 0.32 2.73 $\pm$ 0.35 2.23 $\pm$ 0.27 2.67 $\pm$ 0.35 2.27 $\pm$ 0.33
EchoMimic [8] 2.67 $\pm$ 0.37 3.07 $\pm$ 0.30 2.20 $\pm$ 0.34 2.50 $\pm$ 0.37 2.70 $\pm$ 0.36
FLOAT (Ours) 3.93 $\pm$ 0.21 3.57 $\pm$ 0.33 4.13 $\pm$ 0.27 3.77 $\pm$ 0.30 3.87 $\pm$ 0.30
In Tab.˜7, we conduct a mean opinion score (MOS) based user study to compare the perceptual quality of each method (e.g., teeth clarity and naturalness of emotion). We generate 6 videos by using the baselines and FLOAT, and ask 15 participants to evaluate each generated video with five evaluation factors in the range of 1 to 5. As shown in Tab.˜7, FLOAT outperforms the baselines.
Figure 14: Example of user study interface. (Left) Test Sheet; (Right) Answer Sheet. Participants were asked to evaluate 5 questions for each video (total 180 videos).
In Fig.˜14, we provide an example of test and answer sheet used of the user study. We asked 15 participants to evaluate five questions for each generated video produced by the baselines and FLOAT. Consequently, each participant scores total 180 questions, with responses ranged from 1 to 5. Additionally, we include the supplementary videos used in the user study.
C.4 Video Results
We include video results to further illustrate the performance of our method, including emotion redirection, additional driving conditions, and OOD results. Please refer to provided videos.
Appendix D Discussion
Ethical Consideration   This work aims to advance virtual avatar generation. However, as it can generate realistic talking portrait only from a single image and audio, we considerably recognize the potential for misuse, such as deepfake creation. Attaching watermarks to generated videos and carefully restricted license can mitigate this issues. Additionally, we encourage researchers in deepfake detection to use our results as data to improve detection tools.
Limitation and Further Work   While our method can generate realistic talking portrait video from a single source image and a driving audio, it has several limitations.
First, our method cannot generate more vivid and naunced emotional talking motion. This is because the speech-driven emotion labels are restricted to seven basic emotions, making it challenging to capture more nuanced emotions like shyness. We believe this limitation can be addressed by incorporating textual cues (e.g., “gazing forward with a shyness"), an idea we plan to explore in future work. Moreover, any other approaches to enhance the naturalness of talking motion are key directions for our future work.
Figure 15: Distribution yaw angles in training dataset [97, 46] for FLOAT.
Figure 16: Failure case of FLOAT. It often struggles to handle non-frontal faces and accessories, such as glasses. Please refer to supplementary video.
[…]
This limitations can be mitigated by introducing carefully curated external data, as demonstrated by other concurrent methods [76, 89, 31, 90, 25], or by incorporating multi-view supervision [77] when training our motion latent auto-encoder. We provide examples of failure case in Fig.˜16 and supplementary video.
Acknowledgment   The source images and audio used in this paper are taken from other talking portrait generation methods [76, 90, 89, 8, 96]. We sincerely thank the authors of these works for their valuable contributions. Note that the individuals depicted in our source images and the speech generated in our experiments are not associated with the actual persons they represent.
Figure 17: Detailed Model architecture of our motion latent auto-encoder. The notations are adopted from LIA [85] and StyleGAN2 [33].
Figure 18: Detailed model architecture for constructing the driving conditions $\mathbf{c}_{t}\in\mathbb{R}^{(L^{\prime}+L)\times h}$ in FLOAT.
Figure 19: Out-of-distribution results. The first row shows the result for Chinese audio, and the second row shows the result for singing audio. Please refer to supplementary video.
Figure 20: Ablation on emotion guidance scale $\gamma_{e}$. The predicted speech-to-emotion label is disgust of $99\%9$. Please refer to supplementary video.
Figure 21: Ablation results on frame-wise AdaLN and flow matching. Please refer to supplementary video.
Figure 22: Ablation results on frame-wise AdaLN and flow matching. Please refer to supplementary video.
Figure 23: Ablation results on frame-wise AdaLN and flow matching. Please refer to supplementary video.
Figure 24: Qualitative comparison results with state-of-the-art methods. Please refer to supplementary video.
Figure 25: Qualitative comparison results with state-of-the-art methods. Please refer to supplementary video.
Figure 26: Qualitative comparison results with state-of-the-art methods. Please refer to supplementary video.

### v5

@@
 It can enhance the emotion-related talking motion by leveraging speech-driven emotion labels, a natural way of emotion-aware motion control.
-$\star$\$$\star$\$footnotetext: This work was done during South Korea Mandatory Military Service at DeepBrain AI Inc.
+[…]
 1 Introduction
@@
 We encode the source image $S\in\mathbb{R}^{3\times H\times W}$ into the latent with the explicit identity-motion decomposition $w_{s}=w_{s\to r}+w_{r\to s}\in\mathbb{R}^{d}$.
-Given audio segments $a^{-L^{\prime}:L}\in\mathbb{R}^{(L^{\prime}+L)\times d_{a}}$ of the length $L^{\prime}+L$ and the reference motion $w_{r\to s}$in $\mathbb{R}^{d}$, and the speech-driven emotion label $w_{e}$in $\mathbb{R}^{7}$, a flow matching transformer estimates the generating vector field $v_{t}(\varphi_{t}(x_{0}),\mathbf{c}_{t};\theta)\in\mathbb{R}^{L\times d}$ from noisy motion latents, which is used to solve corresponding ODE and generates the motion latents $w_{r\to\hat{D}^{1:L}}$.
+Given audio segments $a^{-L^{\prime}:L}\in\mathbb{R}^{(L^{\prime}+L)\times d_{a}}$ of the length $L^{\prime}+L$ and the reference motion $w_{r\to s}$ $\in$ $\mathbb{R}^{d}$, and the speech-driven emotion label $w_{e}$ $\in$ $\mathbb{R}^{7}$, a flow matching transformer estimates the generating vector field $v_{t}(\varphi_{t}(x_{0}),\mathbf{c}_{t};\theta)\in\mathbb{R}^{L\times d}$ from noisy motion latents, which is used to solve corresponding ODE and generates the motion latents $w_{r\to\hat{D}^{1:L}}$.
 Finally, the sequence of latents $w_{S\to\hat{D}^{1:L}}:=(w_{S\to r}+w_{r\to\hat{D}^{l}})_{l=1}^{L}$ are decoded into the video $\hat{D}^{1:L}\in\mathbb{R}^{L\times 3\times H\times W}$.
-Let $x$in $\mathbb{R}^{d}$ be a data, $t$in [0,1][0,1][ 0 , 1 ] be the time, and $q$ be a unknown target distribution.
+Let $x$ $\in$ $\mathbb{R}^{d}$ be a data, $t$ $\in$ $[0,1]$ be the time, and $q$ be a unknown target distribution.
 We can define a flow as a time-dependent transformation $\varphi_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that transforms a tractable prior distribution $p_{0}$ to the distribution $p_{1}\approx q$.
@@
 $\mathcal{L}_{\text{FM}}(\theta):=\|v_{t}(x;\theta)-u_{t}(x)\|_{2}^{2},$ (3)
-where $t$sim $\mathcal{U}[0,1]$ and $x$sim $p_{t}(x)$.
+where $t$ $\sim$ $\mathcal{U}[0,1]$ and $x$ $\sim$ $p_{t}(x)$.
 However, the target generating vector field $u_{t}$ and the sample distribution $p_{t}$ are intractable.
-To address this issue, [42] proposes a method for constructing a “conditional" probability path $p_{t}(\cdot|x_{1})$ as well as target “conditional" vector field $u_{t}(\cdot|x_{1})$ using a sample $x_{1}$sim $q$ as a condition.
+To address this issue, [42] proposes a method for constructing a “conditional" probability path $p_{t}(\cdot|x_{1})$ as well as target “conditional" vector field $u_{t}(\cdot|x_{1})$ using a sample $x_{1}$ $\sim$ $q$ as a condition.
 And they prove that the following objective
 $\mathcal{L}_{\text{CFM}}(\theta):=\|v_{t}(x;\theta)-u_{t}(x|x_{1})\|_{2}^{2},$ (4)
-where $t$sim $\mathcal{U}[0,1]$ and $x$sim $p_{t}(x|x_{1})$, is equivalent to (3) with respect to the gradient $\nabla_{\theta}$.
-One natural way of constructing $u_{t}(\cdot|x_{1})$ is a “straight line" that connects $x_{0}$sim $p_{0}$ and $x_{1}$sim $q$, drawing an optimal transport (OT) path with constant velocity [42].
+where $t$ $\sim$ $\mathcal{U}[0,1]$ and $x$ $\sim$ $p_{t}(x|x_{1})$, is equivalent to (3) with respect to the gradient $\nabla_{\theta}$.
+One natural way of constructing $u_{t}(\cdot|x_{1})$ is a “straight line" that connects $x_{0}$ $\sim$ $p_{0}$ and $x_{1}$ $\sim$ $q$, drawing an optimal transport (OT) path with constant velocity [42].
 Specifically, a linear time interpolation between $x_{0}$ and $x_{1}$ gives us the flow $x_{t}=\varphi_{t}(x)=(1-t)x_{0}+tx_{1}$, the conditional probability path $p_{t}(x|x_{1})$ defined via the affine transformation $p_{t}(x|x_{1})=\mathcal{N}(x|tx_{1},(1-t)^{2}I)$, and the target generating vector field $u_{t}(x|x_{1})=x_{1}-x_{0}$.
@@
 $\mathcal{L}_{\text{OT}}(\theta):=\|v_{t}((1-t)x_{0}+tx_{1};\theta)-(x_{1}-x_{0})\|_{2}^{2},$ (5)
-where $t$sim $\mathcal{U}[0,1]$, $x_{0}$sim $p_{0}$, and $x_{1}$sim $q$, all of which are tractable.
+where $t$ $\sim$ $\mathcal{U}[0,1]$, $x_{0}$ $\sim$ $p_{0}$, and $x_{1}$ $\sim$ $q$, all of which are tractable.
 Classifier-free Vector Field   [11] formulates a classifier-free vector field (CFV) technique for flow matching, which enables class-conditional sampling more controllable manner without any extra classifier trained on noisy trajectory.
@@
 We provide an overview of FLOAT in Fig.˜2.
-Given source image $S$in $\mathbb{R}^{3\times H\times W}$, and a driving audio signal $a^{1:L}$in $\mathbb{R}^{L\times d_{a}}$ of length $L$, our method generates a video
+Given source image $S$ $\in$ $\mathbb{R}^{3\times H\times W}$, and a driving audio signal $a^{1:L}$ $\in$ $\mathbb{R}^{L\times d_{a}}$ of length $L$, our method generates a video
 $\hat{D}^{1:L}=(\hat{D}^{l})_{l=1}^{L}\in\mathbb{R}^{L\times 3\times H\times W}$ (7)
@@
 Thus, our first goal for realistic talking portrait is to obtain good motion latent space, capturing both global (e.g., head motion) and fine-grained local (e.g., facial expressions, mouth and pupil movement) dynamics.
-Instead of VAE of SD, we pre-train a motion latent auto-encoder that encodes images into motion latents.
-This is done by training an auto-encoder to reconstruct a driving image from a source image sampled from the same video clip, enforcing the encoder to implicitly capture both temporally adjacent and distant motions.
-Inspired by LIA [85], we introduce a learned orthonormal basis that decomposes the motion along distinct orthogonal directions.
+Instead of VAE of SD, we employ LIA [85] as a base motion latent auto-encoder and pre-train it to encode images into motion latents.
+This is achieved by training the auto-encoder to reconstruct a driving image from a source image sampled from the same video clip, enforcing the encoder to implicitly capture both temporally adjacent and distant motions.
+Following [85], we use a learned orthonormal basis that can decompose the motion along distinct orthogonal directions.
 Specifically, our motion auto-encoder encodes the source $S$ into the latent $w_{S}\in\mathbb{R}^{d}$ with following explicit decomposition:
 $w_{S}:=w_{S\to r}+w_{r\to S},$ (8)
-where $w_{S\to r}$in $\mathbb{R}^{d}$ is the identity latent and
+where $w_{S\to r}$ $\in$ $\mathbb{R}^{d}$ is the identity latent and
 $w_{r\to S}=\sum_{m=1}^{M}\lambda_{m}(S)\cdot\mathbf{v}_{m}\in\mathbb{R}^{d}$ (9)
-is the motion latent with $\lambda(S)$ :=:=:= $\left(\lambda_{m}(S)\right)_{m=1}^{M}$in $\mathbb{R}^{M}$ being the source-dependent motion coefficients that span the learned source-agnostic motion basis $V:=\{\mathbf{v}_{m}\}_{m=1}^{M}\subseteq\mathbb{R}^{d}$.
+is the motion latent with $\lambda(S)$ :=:= $\left(\lambda_{m}(S)\right)_{m=1}^{M}$ $\in$ $\mathbb{R}^{M}$ being the source-dependent motion coefficients that span the learned source-agnostic motion basis $V:=\{\mathbf{v}_{m}\}_{m=1}^{M}\subseteq\mathbb{R}^{d}$.
 In this space, $\lambda_{m}(S)$ is the intensity of the motion direction $\mathbf{v}_{m}$.
@@
 Armed with this linear orthogonal space, we employ OT-based flow matching [42, 44] for the motion sampling.
-Specifically, we predict a vector field $v_{t}(x_{t},\mathbf{c}_{t};\theta)$in $\mathbb{R}^{L\times d}$ where $x_{t}$ is the sample at flow time $t\in[0,1]$, and $\mathbf{c}_{t}$in $\mathbb{R}^{L\times h}$ represents the driving conditions for $L$ consequent frames.
+Specifically, we predict a vector field $v_{t}(x_{t},\mathbf{c}_{t};\theta)$ $\in$ $\mathbb{R}^{L\times d}$ where $x_{t}$ is the sample at flow time $t\in[0,1]$, and $\mathbf{c}_{t}$ $\in$ $\mathbb{R}^{L\times h}$ represents the driving conditions for $L$ consequent frames.
 This vector field generates the flow $\varphi_{t}:[0,1]\times\mathbb{R}^{L\times d}\to\mathbb{R}^{L\times d}$ of $L$ frames by solving ODE (Eq.˜2).
@@
 $\gamma_{i}^{l}\times\text{LN}(X_{t}^{l})+\beta_{i}^{l}\in\mathbb{R}^{h}\quad\text{and}\quad\alpha_{i}^{l}\times X_{t}^{l}\in\mathbb{R}^{h},$ (10)
-[…]
+respectively, where $i$ $\in$ $\{1,2\}$, $h$ is the hidden dimension, $\text{LN}(\cdot)$ denotes layer norm [40], and $X_{t}^{l}$ is the $l$-th input for each operation at flow time $t$ $\in$ $[0,1]$.
+The coefficients $\alpha_{i}^{l},\beta_{i}^{l},\gamma_{i}^{l}\in\mathbb{R}^{h}$ are computed from the condition $\mathbf{c}_{t}^{l}$ $\in$ $\mathbb{R}^{h}$ through a linear layer, ToScaleShift, as depicted in Fig.˜4.
 Speech-driven Emotion Enhancement   How can we make talking motions more expressive and natural?
@@
 Figure 4: Frame-wise vector field predictor block at inference.
-Driving Conditions   We concatenate the audio representation $a^{1:L}$in $\mathbb{R}^{L\times d_{a}}$ of a pre-trained Wav2Vec2.0 [1], the speech emotion label $w_{e}$in $\mathbb{R}^{7}$, and the source motion latent $w_{r\to S}$in $\mathbb{R}^{d}$.
-Next, we add the flow time step embedding $\text{Emb}(t)$in $\mathbb{R}^{h}$ to these conditions, producing $\mathbf{c}_{t}$in $\mathbb{R}^{L\times h}$ via a linear layer, ToCondition, as depicted in Fig.˜2, where $\text{Emb}(t)$ is computed using the sinusoidal position embedding [79].
+Driving Conditions   We concatenate the audio representation $a^{1:L}$ $\in$ $\mathbb{R}^{L\times d_{a}}$ of a pre-trained Wav2Vec2.0 [1], the speech emotion label $w_{e}$ $\in$ $\mathbb{R}^{7}$, and the source motion latent $w_{r\to S}$ $\in$ $\mathbb{R}^{d}$.
+Next, we add the flow time step embedding $\text{Emb}(t)$ $\in$ $\mathbb{R}^{h}$ to these conditions, producing $\mathbf{c}_{t}$ $\in$ $\mathbb{R}^{L\times h}$ via a linear layer, ToCondition, as depicted in Fig.˜2, where $\text{Emb}(t)$ is computed using the sinusoidal position embedding [79].
 Training   We train FLOAT by reconstructing a target vector field computed from driving frames using the corresponding audio segments and a source motion latent.
-We choose a pair of driving motions and corresponding audio ($w_{r\to D^{1:L}}$, $a^{1:L}$), and construct the target vector field $u_{t}(x|w_{r\to D^{1:L}})=w_{r\to D^{1:L}}-x_{0}$in $\mathbb{R}^{L\times d}$ with noisy input $\varphi_{t}(x_{0})=(1-t)x_{0}+tw_{r\to D^{1:L}}$ ($t$sim $\mathcal{U}[0,1]$ and $x_{0}$sim $\mathcal{N}(0^{1:L},I)$).
+We choose a pair of driving motions and corresponding audio ($w_{r\to D^{1:L}}$, $a^{1:L}$), and construct the target vector field $u_{t}(x|w_{r\to D^{1:L}})=w_{r\to D^{1:L}}-x_{0}$ $\in$ $\mathbb{R}^{L\times d}$ with noisy input $\varphi_{t}(x_{0})=(1-t)x_{0}+tw_{r\to D^{1:L}}$ ($t$ $\sim$ $\mathcal{U}[0,1]$ and $x_{0}$ $\sim$ $\mathcal{N}(0^{1:L},I)$).
 For smooth transitions of sequences longer than the window length $L$, we incorporate last $L^{\prime}$ audio features and motion latents $w_{r\to D^{-L^{\prime}:0}}$ from the preceding window as additional input.
@@
 $+\|v_{t}^{-L^{\prime}:0}(x_{t},\mathbf{c}_{t};\theta)-w_{r\to D^{-L^{\prime}:0}}\|,$
-where $x_{t}$ :=:=:= $[w_{r\to D^{-L^{\prime}:0}}|~\varphi_{t}(x_{0})]$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$ is the concatenated input, $\mathbf{c}_{t}$in $\mathbb{R}^{(-L^{\prime}+L)\times h}$ is the driving condition consisting of $[t,w_{r\to S},w_{e},a^{1:L},a^{-L^{\prime}:0}]$.
+where $x_{t}$ :=:= $w_{r\to D^{-L^{\prime}:0}}|~\varphi_{t}(x_{0})] \in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$ is the concatenated input, $\mathbf{c}_{t}$ $\in$ $\mathbb{R}^{(-L^{\prime}+L)\times h}$ is the driving condition consisting of $t,w_{r\to S},w_{e},a^{1:L},a^{-L^{\prime}:0}]$.
 Note that $w_{e}$ and $w_{r\to S}$ are shared across the $L^{\prime}+L$ frames.
@@
 $\mathcal{L}_{\text{vel}}(\theta)=\|\Delta v_{t}-\Delta u_{t}\|,$ (12)
-where $\Delta v_{t}$ and $\Delta u_{t}$ are the one-frame difference along the time-axis for the prediction $v_{t}$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$ and the target $[w_{r\to D^{-L^{\prime}:0}}|~u_{t}]$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$, respectively.
+where $\Delta v_{t}$ and $\Delta u_{t}$ are the one-frame difference along the time-axis for the prediction $v_{t}$ $\in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$ and the target $w_{r\to D^{-L^{\prime}:0}}|~u_{t}] \in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$, respectively.
 The total objective $\mathcal{L}_{\text{total}}(\theta)$ is
@@
 where $\lambda_{\text{OT}}$ and $\lambda_{\text{vel}}$ are the balancing coefficients.
-During training, we apply dropout to $w_{r}$, $w_{e}$, and $a^{1:L}$ with a probability of 0.10.10.1 for CFV.
-Additionally, we apply dropout to the preceding audio and motion latents with a probability 0.50.50.5 for smooth transition in the initial window.
+During training, we apply dropout to $w_{r}$, $w_{e}$, and $a^{1:L}$ with a probability of $0.1$ for CFV.
+Additionally, we apply dropout to the preceding audio and motion latents with a probability $0.5$ for smooth transition in the initial window.
 Table 1: Quantitative comparison results with state-of-the-art methods on HDTF [97] / RAVDESS [46].
 The best result for each metric is in bold, and the second-best result is underlined.
-: evaluated with raw $\times 256$ 256 resolution outputs.
+: evaluated with raw $256\times 256$ resolution outputs.
 Method Image & Video Generation Lip Synchronization
@@
 5.2 Implementation Details
-[…]
+The motion latent dimension is set to $d$ == $512$ with $M$ == 2020 distinct orthogonal directions.
+For the vector predictor, we use 8 attention heads, a hidden dimension $h$ == $1024$, and an attention window length $T$ == 22.
+Considering the length of the training video clips, we set $L$ == $50$ frames with preceding $L^{\prime}$ == $10$ frames at once, encompassing 2.4 seconds of video.
+We employ the Adam optimizer [36] with a batch size of 8 and a learning late of $10^{-5}$.
+We use $L1$ distance for the norm $\|\cdot\|$ in the training objective.
+We set the balancing coefficients to $\lambda_{\text{OT}}$ == $\lambda_{\text{vel}}$ == 11.
+The entire training takes about $2$ days for 2,0002,000k steps on a single NVIDIA A100 GPU.
+We use Euler method [42] for the ODE solver.
 5.3 Evaluation
@@
 5.4 Applications
-Test-time Pose Editing via Orthonormal Basis $V$   Since FLOAT learns the underlying motion latent structure, it is natural to assume that for any sampled motion latent $w_{r\to\hat{D}}$, there exist motion coefficients $\{\lambda_{m}(\hat{D})\}_{m=1}^{M}$ satisfying the representation in Eq.˜9: $w_{r\to\hat{D}}$ $=$ $\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m}$.
+Test-time Pose Editing via Orthonormal Basis $V$   Since FLOAT learns the underlying motion latent structure, it is natural to assume that for any sampled motion latent $w_{r\to\hat{D}}$, there exist motion coefficients $\{\lambda_{m}(\hat{D})\}_{m=1}^{M}$ satisfying the representation in Eq.˜9: $w_{r\to\hat{D}}$ == $\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m}$.
 We can always compute these coefficients in closed form by taking inner products between the sampled motion $w_{r\to\hat{D}}$ and the learned orthonormal basis $V$:
-$\langle w_{r\to\hat{D}},~\mathbf{v}_{k}\rangle=\langle\sum_{m=1}^{M}\lambda_{m}(\hat{D})\cdot\mathbf{v}_{m},~\mathbf{v}_{k}\rangle=\lambda_{k}(\hat{D}),$ (15)
+[…]
 where $\langle\mathbf{v}_{m},\mathbf{v}_{k}\rangle=\delta_{m,k}$ and $\delta$ is Kronecker delta.
@@
 Additional Driving Signals   In Fig.˜7 and Tab.˜2, we experiment with additional driving conditions, head poses and image-driven emotion labels, to explore additional controllability in our method.
-We employ 3DMM head pose parameters $p$in $\mathbb{R}^{6}$ [2] extracted by [13].
-We concatenate a sequence of pose parameters $p^{1:L}$in $\mathbb{R}^{L\times 6}$ with the other driving conditions, and then map them to $c_{t}^{1:L}$in $\mathbb{R}^{L\times h}$.
+We employ 3DMM head pose parameters $p$ $\in$ $\mathbb{R}^{6}$ [2] extracted by [13].
+We concatenate a sequence of pose parameters $p^{1:L}$ $\in$ $\mathbb{R}^{L\times 6}$ with the other driving conditions, and then map them to $c_{t}^{1:L}$ $\in$ $\mathbb{R}^{L\times h}$.
 We also experiment on image-driven emotion [63] for frame-wise emotion control rather than the long-term emotion enhancement.
@@
 The best result for each metric is in bold.
-: Results generated by official implementation ($\times 256$ 256)
+: Results generated by official implementation ($256\times 256$)
 Method FID $\downarrow$ FVD $\downarrow$ LPIPS $\downarrow$ E-FID $\downarrow$ P-FID $\downarrow$
@@
 We incorporate 2-scale discriminator $\mathcal{L}_{\text{full-adv}}$ with the non-saturating loss:
-$\mathcal{L}_{\text{full-adv}}=-\log[\text{Disc}_{\text{full}}(\hat{D})],$ (19)
+[…]
 where Disc denotes a discriminator adopted from [33].
@@
   • It is worth noting that we compare with two superior works EMO [76] and VASA-1 [90] based on their demonstration videos due to their unavailable implementation.
-We highly recommend referring to ‘01_\__EMO_\__VASA-1_\__Comparison///xxxx.mp4’.
+We highly recommend referring to ‘01_\_EMO_\_VASA-1_\_Comparison//xxxx.mp4’.
 B.3 More on Experiments
@@
 Ablation on NFE   In general, increasing the number of function evaluation (NFE) reduces the solution error of ODEs.
-As shown in Tab.˜5, even with small NFE =2=2= 2, FLOAT can achieve competitive image quality (FID) and lip synchronization (LSE-D).
+As shown in Tab.˜5, even with small NFE =2=2, FLOAT can achieve competitive image quality (FID) and lip synchronization (LSE-D).
 However, it struggles to capture consistent and expressive motions (FVD and E-FID), resulting in shaky head motion and a static expression.
@@
 In Fig.˜20, we visualize the effect of different emotion guidance scale $\gamma_{e}$.
-For this experiments, the predicted speech-to-emotion label is disgust with $\%9$ probability.
+For this experiments, the predicted speech-to-emotion label is disgust with $99\%$ probability.
 Notably, as increasing $\gamma_{e}$ from 0 to 2, we can observe that emotion-related expressions and motions are enhanced.
 Figure 13: Ablation results on frame-wise AdaLN and flow matching. Please refer to supplementary video for notable differences.
-[…]
+Ablation on AdaLN and Flow Matching  We conduct ablation study on frame-wise AdaLN by comparing it with a cross-attention.
+We adopt the stand cross-attention mechanism described in [19, 71], using transformer encoder architecture for non-autoregressive sequence modeling.
+We use the same attention mask used in the frame-wise AdaLN, which attends to additional $2T$ adjacent frames for the $l$-th input latent: [l2,l1,l,l+1,l+2][l-2,l-1,l,l+1,l+2].
 To compare against flow matching, we implement two diffusion models with distinct parameterizations: $\epsilon$-prediction and $x_{0}$-prediction.
@@
 $\mathcal{L}_{\text{simple, noise}}(\theta)=\|s(x_{t},\mathbf{c}_{t};\theta)-\epsilon\|_{2}^{2},$ (23)
-where $t$sim $\mathcal{U}[0,1]$, $\epsilon$sim $\mathcal{N}(0^{-L^{\prime}:L},I)$, and the noise input $x_{t}$in $\mathbb{R}^{(L^{\prime}+L)\times d}$ is sampled from a forward diffusion process $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ [27].
-In our case, $x_{t}$ is noisy motion latents at diffusion time step $t$, starting from $t=0$ with $x_{0}$ $=$ $w_{r\to D^{1:L}}$in $\mathbb{R}^{(-L^{\prime}+L)\times d}$.
+where $t$ $\sim$ $\mathcal{U}[0,1]$, $\epsilon$ $\sim$ $\mathcal{N}(0^{-L^{\prime}:L},I)$, and the noise input $x_{t}$ $\in$ $\mathbb{R}^{(L^{\prime}+L)\times d}$ is sampled from a forward diffusion process $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ [27].
+In our case, $x_{t}$ is noisy motion latents at diffusion time step $t$, starting from $t=0$ with $x_{0}$ == $w_{r\to D^{1:L}}$ $\in$ $\mathbb{R}^{(-L^{\prime}+L)\times d}$.
 For $x_{0}$-prediction, we predict a clean sample $x_{0}$, instead of noise [59], by the predictor $s(\cdot;\theta)$ with the following simple loss:
@@
 C.3 User Study
-[…]
+Table 7: Mean opinion score (MOS) study results with $95\%$ confidence interval. The score ranges in 1 to 5. The best result for each metric is in bold.
 Method
@@
 Figure 19: Out-of-distribution results. The first row shows the result for Chinese audio, and the second row shows the result for singing audio. Please refer to supplementary video.
-Figure 20: Ablation on emotion guidance scale $\gamma_{e}$. The predicted speech-to-emotion label is disgust of $99\%9$. Please refer to supplementary video.
+Figure 20: Ablation on emotion guidance scale $\gamma_{e}$. The predicted speech-to-emotion label is disgust of $99.99\%$. Please refer to supplementary video.
 Figure 21: Ablation results on frame-wise AdaLN and flow matching. Please refer to supplementary video.